# Learning the Universe: Learning to Optimize Cosmic Initial Conditions with Non-Differentiable Structure Formation Models

### v1

@@ 1 Introduction
Instead of directly modelling the observed galaxy distribution, one can reframe the task as a statistical problem of inferring the initial conditions of the Universe (Jasche & Wandelt, 2013; Kitaura, 2013; Wang et al., 2014). By incorporating physical models of non-linear gravitational structure formation into a Bayesian framework, this approach establishes a direct link between the observed non-linear cosmic structures and the initial density fields from which they evolved (Jasche & Lavaux, 2019). The interest in initial conditions also stems from the crucial role that the initial conditions, along with the laws of physics, play in determining the dynamical evolution of the Universe and the formation of large-scale structures, including galaxies and galaxy clusters. Accurately inferring the initial conditions is thus essential for interpreting observational data, testing cosmological theories, and probing the nature of dark matter and dark energy.
In the context of field-level inference or reconstruction of cosmic initial conditions, several methods have emerged to handle the complex and high-dimensional datasets provided by next-generation cosmological surveys, including Bayesian forward modelling (Jasche & Wandelt, 2013; Kitaura, 2013; Wang et al., 2014; Ata et al., 2014; Jasche et al., 2015; Jasche & Lavaux, 2017, 2019; Schmidt et al., 2019; Lavaux et al., 2019; Kitaura et al., 2021; Bos et al., 2019; Ata et al., 2021; Doeser et al., 2024), maximum-a-posteriori reconstructions (Seljak et al., 2017; Modi et al., 2018; Feng et al., 2018; Horowitz et al., 2019; Modi et al., 2021a; Horowitz & Melchior, 2023), and machine-learning enhanced approaches (Modi et al., 2018; Modi et al., 2021a; Chen et al., 2023; List et al., 2023; Legin et al., 2023; Shallue & Eisenstein, 2023; Jindal et al., 2023; Bayer et al., 2023a; Doeser et al., 2024; Savchenko et al., 2024, 2025). Field-level inference of initial conditions has also shown promise in joint inference with cosmological parameters (Ramanah et al., 2019; Kostić et al., 2023b; Porqueres et al., 2023; Chen et al., 2024; Andrews et al., 2023, 2024; Nguyen et al., 2024; Krause et al., 2024).
@@
A key challenge in accurately modelling observations lies in capturing the connection between matter and haloes or galaxies, commonly referred to as galaxy bias (e.g. Desjacques et al., 2018). For example, Bartlett et al. (2024) showed that non-linear but local galaxy models are insufficient. As high-dimensional field-level inference methods incorporating non-linear structure formation and perturbative or phenomenological galaxy bias models have now become computationally feasible, the next pressing challenge involves how to integrate non-differentiable models in the data analysis. These include halo population models, galaxy formation models, and hydrodynamical simulations, which offer high-fidelity modelling of, e.g. baryonic physics, massive neutrinos, and feedback-reactions (see, e.g. Villaescusa-Navarro et al., 2021; Schaye et al., 2023; Pakmor et al., 2023). Currently, no differentiable algorithms exist for these models, and the implementation and testing of these methods for science-ready applications will require significant development effort and resources to be invested by the community. This is also true for efforts to circumvent traditional non-differentiable models, such as fast differentiable bias models (Modi et al., 2018; Charnock et al., 2020; Ding et al., 2024; Pandey et al., 2024a, b) to replace algorithms such as spherical overdensity halo finders AHF (Knollmann & Knebe, 2009) and ROCKSTAR (Behroozi et al., 2013). Similar efforts also aim to directly learn the mapping between dark matter and galaxy fields through high-fidelity hydrodynamic simulations (Villaescusa-Navarro et al., 2021; Sether et al., 2024; Bourdin et al., 2024; Horowitz & Lukić, 2025).
@@ 3 Method: Gradient-free optimization
Our goal is to address the inverse problem of reconstructing the Universe’s initial conditions from volumetric datasets, such as dark matter halo or galaxy count data. This task poses significant challenges due to the inclusion of complex, non-linear, and often non-differentiable cosmological simulations to casually connect the initial conditions with data. For this work, we explore handling non-differentiability by seeking a maximum likelihood estimate.
@@ 3.2 Incorporation of ML to learn black-box optimization
We seek to learn an alternative to the vanilla, and hand-crafted, gradient-based optimization scheme (see e.g. review by Ruder, 2016), which follows
@@
where $\gamma\in\mathbb{R}_{+}$ is the learning rate. For the data discrepancy we choose a voxel-wise mean-squared likelihood
$\mathcal{L}(\boldsymbol{x}_{t})=\frac{1}{2}(\boldsymbol{d}-\mathscr{S}(\boldsymbol{x}_{t}))^{\top}\Sigma(\boldsymbol{d}-\mathscr{S}(\boldsymbol{x}_{t})),$ (5)
with $\Sigma$ chosen to an identity covariance matrix. As $\nabla_{\boldsymbol{x}_{t}}f(\boldsymbol{x}_{t})\propto\nabla_{\boldsymbol{x}_{t}}\mathcal{L}(\boldsymbol{x}_{t})\propto\nabla_{\boldsymbol{x}_{t}}\mathscr{S}(\boldsymbol{x}_{t})$ and $\nabla_{\boldsymbol{x}_{t}}\mathscr{S}(\boldsymbol{x}_{t})$ is not accessible, we introduce a machine learning neural network (NN) model to learn the mapping from the data discrepancy $\Delta\boldsymbol{d}_{t}$ to the update direction $\Delta\boldsymbol{x}_{t}$ for the initial conditions. In particular, we then have
$\nabla_{\boldsymbol{x}_{t}}\mathcal{L}(\boldsymbol{x}_{t})\propto\underbrace{\underbrace{\left(\boldsymbol{d}-\mathscr{S}(\boldsymbol{x}_{t})\right)}_{\textstyle\Delta\boldsymbol{d}_{t}}\nabla_{\boldsymbol{x}_{t}}\mathscr{S}(\boldsymbol{x}_{t})}_{\textstyle\mathrm{NN}(\Delta\boldsymbol{d}_{t})=\Delta\boldsymbol{x}_{t}},$ (6)
where $\Delta\boldsymbol{d}_{t}$ is computable, while $\nabla_{\boldsymbol{x}_{t}}\mathscr{S}(\boldsymbol{x}_{t})$ is not. Experiments on including $\boldsymbol{x}_{t}$ as an additional input to NN did not show any improvement and we leave further investigation to future studies. We also note there can be $n_{d}$ quantities in the data space that are informative about the initial state. Thus, in general, we have
@@
We also leverage the exploration-exploitation philosophy central to evolutionary computation techniques by incorporating two competing terms: one deterministic and one stochastic (Holland, 1975). Similar to e.g. Fornasier et al. (2021); Riedl et al. (2023) we introduce the latter by a new term $f(\boldsymbol{x}_{t})\mathcal{N}_{t}$ to Eq. (4) for some function $f(\boldsymbol{x}_{t})$ that vanishes as we approach the ground truth and a Gaussian random vector $\mathcal{N}_{t}$ with zero mean and unit diagonal covariance matrix. We choose an $\boldsymbol{x}$-independent stochastic function $f(\boldsymbol{x}_{t})=0.1\times 0.9^{t}$. While the deterministic term from the machine-learning optimizer guides the optimization towards promising regions, the stochastic term introduces a diffusion mechanism that injects randomness into the dynamics to increase the exploration across the search space. The combined update and query rule can now be defined as:
$\boldsymbol{x}_{t+1}$ $=\boldsymbol{x}_{t}-\gamma_{t}\Delta\boldsymbol{x}_{t}=\boldsymbol{x}_{t}-\gamma_{t}[\mathrm{NN}(\Delta\boldsymbol{d}_{t})+f(\boldsymbol{x}_{t})\mathcal{N}_{t}]$ (8)
@@ 3.3.2 Neural loss function
$L_{\mathrm{csl}}=1-\cos(\Delta\boldsymbol{x}_{\mathrm{true}},\Delta\boldsymbol{x}_{\mathrm{pred}})=1-\frac{\Delta\boldsymbol{x}_{\mathrm{true}}^{\top}\Delta\boldsymbol{x}_{\mathrm{pred}}}{||\Delta\boldsymbol{x}_{\mathrm{true}}||\cdot||\Delta\boldsymbol{x}_{\mathrm{pred}}||},$ (10)
@@ 4.2 Non-linear, non-differentiable cosmological simulator
To translate between the initial conditions $\boldsymbol{x}$ and the data space, we employ a high-fidelity $N$-body simulation coupled with a non-differentiable spherical halo finder. The simulations are performed with $128^{3}$ particles within a cubic volume with a side length of $250h^{-1}$ Mpc, i.e. at a Lagrangian resolution of $1.95h^{-1}$ Mpc. The resulting particles and haloes are discretized onto grids of $128^{3}$ elements using a cloud-in-cell mass assignment scheme to obtain the density contrasts. The dark matter overdensity and halo overdensity thus maintain a resolution of $1.95h^{-1}$ Mpc, which enables accurate representation of non-linear cosmic structures.
@@ 4.2.1 $N$-body simulation
$N$-body simulations are currently the most advanced numerical technique for modeling the full non-linear structure formation of the Universe (Vogelsberger et al., 2020; Angulo & Hahn, 2022). In this work, we use the $N$-body simulation code Gadget-IV (Springel et al., 2021). From the white-noise field, the simulator initial conditions of particles are generated at $z=127$ using MUSIC (Hahn & Abel, 2011) together with transfer function from Eisenstein & Hu (Eisenstein & Hu, 1998, 1999). The cosmological parameters used are $\Omega_{\mathrm{m}}=0.3175$, $\Omega_{\mathrm{b}}=0.049$, $h=0.6711$, $n_{\mathrm{s}}=0.9624$, $\sigma_{8}=0.834$, and $w=-1$ consistent with latest constraints by Planck (Aghanim et al., 2020) and equivalent with the fiducial cosmology of the Quijote simulations (Villaescusa-Navarro et al., 2020). We compile and run the dark-matter-only and TreePM code version of Gadget-IV with $\mathrm{PMGRID}=256$. Although in theory differentiable, achieving this would require storing the full phase-space information at every time step of the simulator integrator—on the order of thousands—which memory-wise is infeasible.
@@ 4.3.1 Size of training data
[…]
@@ 4.3.2 Training set-up
Parallel processing of large volumes into sub-boxes is enabled. In particular, we crop the $128^{3}$ fields to $64^{3}$ due to memory requirements and use a batch size of $16$ over multiple NVIDIA A100100100100 $40$GB Tensor Core GPUs. Through periodic padding of the input to the network by $48$ voxels on each side and the rotation and flip augmentations our implementation preserves both translation and (approximately) rotational symmetry. We use an initial learning rate of $4\times 10^{-4}$ with the AdamW optimiser (Loshchilov & Hutter, 2019), which is an extension to the Adam optimiser (Kingma & Ba, 2014) using decoupled weight decay. We use a decay coefficient $\lambda=1\times 10^{-4}$ and parameters $\beta_{1}=0.9$ and $\beta_{2}=0.99$. At a plateau, if the loss is not reduced by a factor $0.999$ over $3$ epochs, we half the learning rate. To achieve high performance, we primarily monitor the cosine similarity loss for the training and validation sets, but the performance after incorporation within the full optimization process is also evaluated.
@@ 4.3.3 Performance
As an initial assessment of the trained model’s performance, we evaluate it on the validation set, while acknowledging that the true evaluation will occur on the test set, that is when the neural optimizer is applied in LULO to reconstruct the initial conditions. The cosine similarity ($1-\mathrm{loss}$) achieved on the validation set is $0.087\pm 0.024$. Applying a Gaussian smoothing kernel with a scale of $10h^{-1}$ Mpc, corresponding to $k\sim 0.3h$ Mpc-1, to both the predicted $\Delta\boldsymbol{x}_{\mathrm{pred}}$ and the true $\Delta\boldsymbol{x}_{\mathrm{true}}$, reveals a higher alignment of cosine similarity $0.944\pm 0.036$ at these larger scales.
@@ 5.1 Initialization
The initial guess for the initial conditions is chosen as a zero field. Given the high-dimensional space of $128^{3}\sim 10^{6}$ voxels and the Gaussian nature of the true initial conditions, characterized by a zero mean and unit variance, a zero field is closer to the truth than a randomly generated white noise field. Due to the reduced initial cosmological power, the first simulation yields no haloes. This means the first input to the neural optimizer is $\Delta\boldsymbol{d}=\boldsymbol{d}$, from which the first search direction $\Delta\boldsymbol{x}$ is predicted. The first mapping initiates the optimization process of iteratively updating the initial conditions per Eqs. (8)–(9) such that they, after forward simulation, increasingly match the data. The full monitoring of the subsequent optimization progress and iterative improvements is presented in Appendix C.

### v2

@@ 1 Introduction
Instead of directly modelling the observed galaxy distribution, the task can be reframed as a statistical problem of inferring the initial conditions of the Universe (Jasche & Wandelt, 2013; Kitaura, 2013; Wang et al., 2013, 2014). By incorporating physical models of non-linear gravitational structure formation into a Bayesian framework, this approach establishes a direct link between the observed non-linear cosmic structures and the initial density fields from which they evolved (see data applications in e.g. Jasche et al., 2015; Lavaux & Jasche, 2016; Jasche & Lavaux, 2019; Kitaura et al., 2021; Lavaux et al., 2019; Ata et al., 2021). Advancing the forward modelling within these Bayesian frameworks (as in e.g. Ata et al., 2014; Schmidt et al., 2019; Jasche & Lavaux, 2019; Bos et al., 2019; Stopyra et al., 2023; Doeser et al., 2024) is essential for effectively handling the complex, high-dimensional data from upcoming cosmological surveys.
In addition to Bayesian forward modelling, several other approaches to infer the initial conditions have been explored, ranging from maximum-a-posteriori (MAP) estimation (Seljak et al., 2017; Feng et al., 2018; Horowitz et al., 2019; Horowitz & Melchior, 2023; Bayer et al., 2023a) to recent machine learning methods. These include recurrent neural networks (Modi et al., 2021a), auto-regressive models (List et al., 2023), score-based diffusion sampling (Legin et al., 2023), final-to-initial mappings (Jindal et al., 2023), neural-enhanced forward modelling (Modi et al., 2018; Doeser et al., 2024), and variational self-boosted sampling (Modi et al., 2023). Convolutional neural networks have further been widely used to improve small-scale reconstructions (e.g. Shallue & Eisenstein, 2023; Chen et al., 2023; Parker et al., 2025; Bottema et al., 2025; Savchenko et al., 2024, 2025).
The interest in initial conditions stems from the crucial role that they, along with the laws of physics, play in determining the dynamical evolution and formation of the large-scale structure of galaxies in the Universe. Accurately inferring the initial conditions is thus essential for interpreting observational data, testing cosmological theories, and probing the nature of dark matter and dark energy. Joint inference of initial conditions and cosmological parameters have shown promise, such as for $(\Omega_{\mathrm{m}},\omega_{0})$ (Ramanah et al., 2019), $\sigma_{8}$ (Kostić et al., 2023b; Nguyen et al., 2024), $(\Omega_{\mathrm{m}},\sigma_{8})$ (Porqueres et al., 2023; Krause et al., 2024), and primordial non-Gaussianties (Chen et al., 2024; Andrews et al., 2023, 2024; Flöss & Meerburg, 2024).
@@
A key challenge in accurately modelling observations lies in capturing the connection between matter and haloes or galaxies, commonly referred to as galaxy bias (e.g. Desjacques et al., 2018). For example, Bartlett et al. (2024) showed that non-linear but local galaxy models are insufficient. As high-dimensional field-level inference methods incorporating non-linear structure formation and perturbative or phenomenological galaxy bias models have now become computationally feasible, the next pressing challenge involves how to integrate non-differentiable models in the data analysis. These include halo population models, galaxy formation models, and hydrodynamical simulations, which offer high-fidelity modelling of, e.g. baryonic physics, massive neutrinos, and feedback-reactions (see, e.g. Villaescusa-Navarro et al., 2021; Schaye et al., 2023; Pakmor et al., 2023). Currently, no differentiable algorithms exist for these models, and the implementation and testing of these methods for science-ready applications will require significant development effort and resources to be invested by the community. This is also true for efforts to circumvent traditional non-differentiable models, such as fast differentiable bias models (Modi et al., 2018; Kodi Ramanah et al., 2019; Charnock et al., 2020; Ding et al., 2024; Pandey et al., 2024a, b) to replace algorithms such as spherical overdensity halo finders AHF (Knollmann & Knebe, 2009) and ROCKSTAR (Behroozi et al., 2013). Similar efforts also aim to directly learn the mapping between dark matter and galaxy fields through high-fidelity hydrodynamic simulations (Villaescusa-Navarro et al., 2021; Sether et al., 2024; Bourdin et al., 2024; Horowitz & Lukić, 2025).
@@ 3 Method: Gradient-free optimization
Our goal is to address the inverse problem of reconstructing the Universe’s initial conditions from volumetric datasets, such as dark matter halo or galaxy count data. This task poses significant challenges due to the inclusion of complex, non-linear, and often non-differentiable cosmological simulations to causally connect the initial conditions with data. For this work, we explore handling non-differentiability by seeking a maximum likelihood estimate.
@@ 3.2 Incorporation of ML to learn black-box optimization
Gradient descent (see e.g. review by Ruder, 2016) is based on explicit differentiability as
@@
where $\gamma\in\mathbb{R}_{+}$ is the learning rate. As $f(\boldsymbol{x_{t}})=\mathcal{L}(\mathscr{S}(\boldsymbol{x_{t}}))$, the gradient $\nabla_{\boldsymbol{x}_{t}}f(\boldsymbol{x}_{t})$ is given by the chain rule, which involves differentiating through the simulator $\mathscr{S}$. Since $\mathscr{S}$ is non-differentiable, we seek to learn an alternative to this optimization scheme without relying on gradients. For the data discrepancy, we choose a voxel-wise mean-squared likelihood
$\mathcal{L}(\boldsymbol{x}_{t})=\frac{1}{2}(\boldsymbol{d}-\mathscr{S}(\boldsymbol{x}_{t}))^{\top}\Sigma^{-1}(\boldsymbol{d}-\mathscr{S}(\boldsymbol{x}_{t})),$ (5)
with $\Sigma$ chosen to an identity covariance matrix. Based on Eq. (4), the update step $\Delta\boldsymbol{x}_{t}=\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}$ for the $i$:th voxel can be expressed
$\Delta\boldsymbol{x}_{t}]_{i}\propto[\nabla_{\boldsymbol{x}_{t}}f(\boldsymbol{x}_{t})]_{i}=\sum_{j}\underbrace{\left(\boldsymbol{d}_{j}-[\mathscr{S}(\boldsymbol{x}_{t})]_{j}\right)}_{[\Delta\boldsymbol{d}_{t}]_{j}}\frac{\partial[\mathscr{S}(\boldsymbol{x}_{t})]_{j}}{\partial[\boldsymbol{x}_{t}]_{i}}.$ (6)
As $\Delta\boldsymbol{d}_{t}$ can be computed, while the partial derivatives of $\mathscr{S}$ are not available, we introduce a neural network (NN) model to learn the mapping from the data discrepancy $\Delta\boldsymbol{d}_{t}$ to the update step $\Delta\boldsymbol{x}_{t}$ for the initial conditions, i.e., $\Delta\boldsymbol{x}_{t}=\textstyle\mathrm{NN}(\Delta\boldsymbol{d}_{t})$. This bypasses the need to differentiate through the simulator $\mathscr{S}$. Experiments on including $\boldsymbol{x}_{t}$ as an additional input to the NN did not show any improvement, and we leave further investigation to future studies. We also note that there can be $n_{d}$ quantities in the data space that are informative about the initial state. Thus, in general, we have
@@
We leverage the exploration-exploitation philosophy central to evolutionary computation techniques by incorporating two competing terms: one deterministic and one stochastic (Holland, 1975). Similar to e.g. Fornasier et al. (2021); Riedl et al. (2023) we introduce the latter by a new term $B(\boldsymbol{x}_{t})\mathcal{N}_{t}$ to Eq. (4) for some function $B(\boldsymbol{x}_{t})$ that vanishes as we approach the ground truth and a standard Gaussian random vector $\mathcal{N}_{t}$. We choose an $\boldsymbol{x}$-independent function $B(\boldsymbol{x}_{t})=0.1\times 0.9^{t}$ and draw a new $\mathcal{N}_{t}$ at each $t$. While the deterministic term from the machine-learning optimizer guides the optimization towards promising regions, the stochastic term introduces a diffusion mechanism that injects randomness into the dynamics to increase the exploration across the search space. The combined update and query rule can now be defined as:
$\boldsymbol{x}_{t+1}$ $=\boldsymbol{x}_{t}-\gamma_{t}\Delta\boldsymbol{x}_{t}=\boldsymbol{x}_{t}-\gamma_{t}[\mathrm{NN}(\Delta\boldsymbol{d}_{t})+B(\boldsymbol{x}_{t})\mathcal{N}_{t}]$ (8)
@@ 3.3.2 Neural loss function
[…]
@@ 4.2 Non-linear, non-differentiable cosmological simulator
To translate between the initial conditions $\boldsymbol{x}$ and the data space, we employ a high-fidelity $N$-body simulation coupled with a non-differentiable spherical halo finder. The simulations are performed with $128^{3}$ particles within a cubic volume with a side length of $250h^{-1}$ Mpc, i.e. at a Lagrangian resolution of $1.95h^{-1}$ Mpc. The resulting particles and haloes are discretized onto grids of $128^{3}$ elements using a cloud-in-cell mass assignment scheme to obtain the density contrasts. The dark matter overdensity and halo overdensity thus maintain a resolution of $1.95h^{-1}$ Mpc, which enables accurate representation of non-linear cosmic structures. The number density of haloes, used as tracers in our mock setup, is $\sim 2.2\times 10^{-4}$ $(h/\mathrm{Mpc})^{3}$, which is comparable to the tracer number density in galaxy surveys (see e.g. Ata et al., 2017; Collaboration et al., 2024).
@@ 4.2.1 $N$-body simulation
NN-body simulations are currently the most advanced numerical technique for modeling the full non-linear structure formation of the Universe (Vogelsberger et al., 2020; Angulo & Hahn, 2022). In this work, we use the $N$-body simulation code Gadget-IV (Springel et al., 2021). From the white-noise field, the simulator initial conditions of particles are generated at $z=127$ using MUSIC (Hahn & Abel, 2011) together with transfer function from Eisenstein & Hu (Eisenstein & Hu, 1998, 1999). The cosmological parameters used are $\Omega_{\mathrm{m}}=0.3175$, $\Omega_{\mathrm{b}}=0.049$, $h=0.6711$, $n_{\mathrm{s}}=0.9624$, $\sigma_{8}=0.834$, and w=1w=-1 consistent with latest constraints by Planck (Aghanim et al., 2020) and equivalent with the fiducial cosmology of the Quijote simulations (Villaescusa-Navarro et al., 2020). We compile and run the dark-matter-only and TreePM code version of Gadget-IV with $\mathrm{PMGRID}=256$. Although in theory differentiable, achieving this would require storing the full phase-space information at every time step of the simulator integrator—on the order of thousands—which memory-wise is infeasible.
@@ 4.3.1 Size of training data
We generate $1000$ random initial conditions, run the simulation pipeline, and use Eq. (11) to perturb each initial condition $5$ times. For each of the $6000$ simulations, we store the initial white noise field $\boldsymbol{x}$, the redshift $z=0$ snapshot from GADGET-IV as well as the halo catalog provided by AHF. In total, we create $5000$ difference fields for the initial conditions $\Delta\boldsymbol{x}$ as well as for the data $\Delta\boldsymbol{d}=[\Delta\boldsymbol{\delta}^{\mathrm{halo}}_{\mathrm{mw}},\Delta\boldsymbol{\delta}^{\mathrm{halo}}_{\mathrm{count}}]$.
@@ 4.3.2 Training set-up
Parallel processing of large volumes into sub-boxes is enabled. In particular, we crop the $128^{3}$ fields to $64^{3}$ due to memory requirements and use a batch size of $16$ over multiple NVIDIA A100100 4040GB Tensor Core GPUs. We preserve translation symmetry by periodically padding each sub-box by $48$ voxels on each side. Additionally, we randomly reflect and/or apply rotations of the sub-boxes in multiples of $90$ degrees, thus covering a discrete set of orientations rather than all possible rotations, resulting in rotational symmetry being approximately preserved. We use an initial learning rate of $4\times 10^{-4}$ with the AdamW optimiser (Loshchilov & Hutter, 2019), which is an extension to the Adam optimiser (Kingma & Ba, 2014) using decoupled weight decay. We use a decay coefficient $\lambda=1\times 10^{-4}$ and parameters $\beta_{1}=0.9$ and $\beta_{2}=0.99$. At a plateau, if the loss is not reduced by a factor $0.999$ over $3$ epochs, we half the learning rate. To achieve high performance, we primarily monitor the cosine similarity loss for the training and validation sets, but the performance after incorporation within the full optimization process is also evaluated.
@@ 4.3.3 Performance
As an initial assessment of the trained model’s performance, we evaluate it on the validation set, while acknowledging that the true evaluation will occur on the test set, that is when the neural optimizer is applied in LULO to reconstruct the initial conditions. The cosine similarity ($1-\mathrm{loss}$) achieved on the validation set is $0.087\pm 0.024$. Applying a Gaussian smoothing filter with standard deviation $10h^{-1}$ Mpc to both the predicted $\Delta\boldsymbol{x}_{\mathrm{pred}}$ and the true $\Delta\boldsymbol{x}_{\mathrm{true}}$, reveals a higher alignment of cosine similarity $0.944\pm 0.036$ at these larger scales.
@@ 5.1 Initialization
The initial guess for the initial conditions is chosen as a zero field. Given the high-dimensional space of $128^{3}\sim 10^{6}$ voxels and the Gaussian nature of the true initial conditions, characterized by a zero mean and unit variance, a zero field is closer to the truth than a randomly generated white noise field. Due to the reduced initial cosmological power, the first simulation yields no haloes. This means the first input to the neural optimizer is $\Delta\boldsymbol{d}=\boldsymbol{d}$, from which the first search direction $\Delta\boldsymbol{x}$ is predicted. The first mapping initiates the optimization process of iteratively updating the initial conditions per Eqs. (8)–(9) such that they, after forward simulation, increasingly match the data. The full monitoring of the subsequent optimization progress and iterative improvements is presented in Appendix C.1.